The solubility of Ca(OH) 2 ·8H 2 O in water at 25 ◦ C is 1.73 g per litre. What is the formality and molality of a saturated solution at 25 ◦ C, assuming it to be a very dilute solution? [Ca =

40, O = 16 and H = 1]


1
Expert's answer
2021-10-28T07:48:53-0400

Molar mass of Ca(OH)2= 74.093 g/mol

So, number of moles of Ca(OH)

2

in 1 kg of water= 1.73/74.093

= 0.0233mol

1 mole of Ba(OH) 2 produces 2 moles of OH- ions

= 2×0.0233= 0.0466

Hence molality = 0.0466m